I really need help with this stuff I get so confused.
Given f(x)=-4x-4, find f(2)

Answer by blondesrbetr37(2)

What you do is you plug 2 in for x so the problem reads f(2)=-4(2)-4 so f(2)=-8-4 so f(2)=-12 so -12 is your answer!